This is an application problem int he quadratic equation chapter.
Thomas is going to make an open-top box by cutting equal squares from the four corners of an 11 inch by 14 inch sheet of cardboard and folding up the sides. If the area of the base is to be 80 square inches, then what size square should be cut from each corner?
I cannot figure out how to begin on this, so I really appreciate your help.

Answer by scott8148(6628)

the length and width of the sheet will both be reduced by 2 corner square sizes (one at each end)
___ these reduced dimensions are the length and width of the base

(11 - 2x) (14 - 2x) = 80

154 - 50x + 4x^2 = 80

2x^2 - 25x + 37 = 0

use quadratic formula to find x